• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:00:46 on October31,2002

    Entry number 89383

    keyword=Shift Summary

    Day Shift: Halloween 2002
    John LeRose: SL
    Doug Higinbotham: TO
    Paul Ulmer: 3rd (arrives late, car crash in the tunnel. H. Ibrahim stays an extra hour)
    At start of shift Owl shift was just finishing tracking down an IOC problem.
    Software reset fixed it. There was some confusion caused by looking at IOC's as reported by MCC instead of the Hall A IOC's (use the "Hall A IOC Systems" window to diagnose problems)
    8:15 run 2881 started (Q3D_f40l)
    8:50 MCC says they have a vacuum problem in the Compton Chicane
    9:02 MCC making a controlled access to investigate the problem
    10:45 Access completed (ep turbo pump replaced) Called for return to beam permit
    Of course we are now in machine development time.
    11:27 Machine development over MCC says they're ready to give us beam
    11:34 start run 2882
    12:15 discover FPP not working during run 2881. Problem was thresholds were set to zero (probably happened during resets at 0800). We reset them during 2882.
    12:25 Requested MCC to put 1C12 otr in
    12:26 start run 2883
    12:35 MCC having trouble getting beam past the OTR. We take it out.
    13:25 start 2884
    14:27 start 2885
    15:19 start 2886
    2886 Continuing at shift end